# FlowSteer: Interactive Agentic Workflow Orchestration via End-to-End Reinforcement Learning

### v1

@@ 1 Introduction
In recent years, a variety of powerful agentic systems have been applied to solve a wide range of human problems (Ruan et al., 2023; Shen et al., 2023; Hong et al., 2024), gradually moving beyond single-turn question answering (QA) toward executable end-to-end task completion. In this process, workflow orchestration has become a key bridge from task goals to reproducible execution: by organizing operators into an executable workflow graph, systems can complete complex tasks with improved controllability, debuggability, and reusability (Zeng et al., 2023; Qian et al., 2024), as shown in Figure 1. However, in practice, workflow construction still heavily relies on manual drag-and-drop and rule-based configuration (Li et al., 2025), making it costly to transfer across new tasks, new operator libraries, new model backends, or different application domains.
@@
To address these issues, three main paradigms of workflow orchestration have emerged, as shown in Figure 2. First, static workflow selection retrieves pre-defined workflows from a library based on task similarity (Wang et al., 2024). Second, offline workflow generation trains policy models via supervised fine-tuning (SFT) or group relative policy optimization (GRPO) (Shao et al., 2024b) to generate workflows. Third, automated workflow optimization—such as AFlow (Zhang et al., 2024a), GPTSwarm (Zhuge et al., 2024), and LATS (Zhou et al., 2023)—combines search and execution feedback to iteratively improve workflow structures.
@@
However, these methods still face several challenges: (i) High manual/heuristic dependence—rules and templates require continual maintenance and are often tightly coupled to specific scenarios, limiting reuse and generalization (Schick et al., 2023; Wang et al., 2024); (ii) Operator/backend lock-in (path lock-in)—existing approaches tend to rely on fixed operators or a single strong large language model (LLM) backend, making it difficult for a policy model to compose operators in a plug-and-play manner, so performance and robustness drop sharply when the operator library or backend environment changes (Yao et al., 2023; Zhou et al., 2024); (iii) Sparse and unstable learning signals—training with only terminal correctness rewards can lead to shortcut behaviors (e.g., premature termination and oversimplified graphs) and reward hacking, and may further suffer from intra-group advantage collapse, making long-horizon credit assignment unstable (Yu et al., 2024; DeepSeek-AI, 2025).
@@ 2 Related Work
Agent Workflows. Before the rise of LLM agents, workflow automation was largely rule- or template-driven, with operators and control flow specified by hand. In the era of LLM agents, agent workflows improve long-horizon reliability via a plan–act–feedback loop (Huang et al., 2025; Shao et al., 2024a; Hong et al., 2024). Existing studies can be grouped into three lines: single-agent decision making, which models tool use as sequential decisions (Huang et al., 2025) or interleaved reasoning and acting (Tam et al., 2024); orchestration, which uses LLM controllers for tool/model routing (Shao et al., 2024a) or constrained application programming interface (API) planning to ground intent (Wang et al., 2024); and multi-agent collaboration, which relies on standard operating procedures (SOPs)/roles (Hong et al., 2024) or cross-team orchestration (Chen et al., 2025a). For sustained capability, agentic tool reasoning (Yao et al., 2025; Luo et al., 2025c) and reusable workflow memory further improve efficiency (Wang et al., 2024; Luo et al., 2025b), supporting scalable automation in practice (Qian et al., 2024). In this paper, we propose FlowSteer, an end-to-end RL framework that learns workflow orchestration from executable canvas feedback.
Reinforcement Learning for Agents. With recent progress in RL for LLM-based agents, agent RL models interaction as a long-horizon Markov decision process (MDP) (Zhou et al., 2024). In hierarchical multi-turn RL, coupling value learning with token-level policy learning eases delayed credit assignment (Zhou et al., 2024; Shen et al., 2026). For tool-chain control, step-grained shaping (Yu et al., 2024) and outcome feedback improve tool selection and self-correction (Feng et al., 2025). When retrieval is treated as an action, search rollouts learn think-then-search behaviors (Jin et al., 2025; Luo et al., 2025a). Moreover, large-scale RL motivates GRPO-style objectives based on verifiable or group-relative signals (DeepSeek-AI, 2025; Shao et al., 2024b), and recent work explores end-to-end RL for collaborative prompting (Liu et al., 2025) and knowledge graph reasoning (Shen et al., 2025a, b).
@@ 4.3 Outcome-Directed End-to-End Reinforcement Learning
which measures the ratio between current policy $\pi_{\theta}$ and behavior policy $\pi_{\theta_{\mathrm{old}}}$ used for trajectory sampling, and the normalized advantage $\hat{A}(\tau_{i})=(R(\tau_{i})-\mu_{\mathrm{src}})/(\sigma_{\mathrm{src}}+\epsilon)$ uses within-group statistics (Shao et al., 2024b) $\mu_{\mathrm{src}}$, $\sigma_{\mathrm{src}}$ partitioned by data source. The mask $\mathrm{mask}_{t}^{(i)}\in\{0,1\}$ is set to 1 for policy-generated tokens and 0 for environment feedback, ensuring gradients only affect policy decisions; $|\tau_{i}|_{\mathrm{mask}}$ is the masked token count. The $\mathrm{clip}(\cdot)$ operator stabilizes updates (Schulman et al., 2017), and the KL term regularizes toward reference $\pi_{\mathrm{ref}}$ with strength $\beta$.
@@ 5.1 Experimental Setup
Datasets. We evaluate FlowSteer on six in-distribution (IID) benchmarks covering three task categories: (i) mathematical reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b)), (ii) question answering (HotPotQA (Yang et al., 2018), SQuAD v2 (Rajpurkar et al., 2018)), and (iii) code generation (MBPP (Austin et al., 2021), HumanEval (Chen et al., 2021)). To assess generalization, we further test on six out-of-distribution (OOD) benchmarks: TriviaQA (Joshi et al., 2017), NaturalQuestions (Kwiatkowski et al., 2019), MathQA (Amini et al., 2019), AIME 2025, APPS (Hendrycks et al., 2021a), and DS-1000 (Lai et al., 2023). More details of the datasets are provided in Appendix D.
Baselines. We compare FlowSteer against the following baselines: (i) direct LLM baselines (Qwen3-8B (Qwen Team, 2025), GPT-4o-mini (OpenAI, 2024)), (ii) standard fine-tuning methods (SFT, GRPO (Shao et al., 2024b)), (iii) workflow-based methods (AFlow (Zhang et al., 2024a)), and (iv) agent with RL methods (AgentFlow (Chen et al., 2025b), Router-R1 (Yuan et al., 2025), Orchestrator (Yang et al., 2025)). More details are provided in Appendix E.
@@ 5.6 Comparison of RL Algorithms (RQ5)
As shown in Table 6, we compare three RL algorithms (DAPO, GRPO, and CWRPO) on six IID benchmarks under identical training settings. CWRPO achieves optimal results across all task types with more stable training dynamics. The pairwise comparison matrices in Figure 5(c–e) show consistent superiority of CWRPO across math, QA, and code tasks. Unlike DAPO and GRPO that primarily focus on answer correctness, CWRPO jointly optimizes structural diversity and task performance through diversity-constrained rewards with conditional release. This design enables the policy to learn not only how to solve problems but also how to orchestrate workflows effectively, achieving better trade-off between workflow quality and answer accuracy.
@@ Appendix D Dataset Details
$\bullet$ AIME 2025: Problems from the 2025 American Invitational Mathematics Examination, representing challenging competition-level mathematics. AIME problems require creative problem-solving and deep mathematical insight, with answers being integers from 0 to 999.
@@ E.2 Fine-Tuning Baselines
$\bullet$ GRPO (Group Relative Policy Optimization) (Shao et al., 2024c): Group Relative Policy Optimization is a reinforcement learning algorithm that normalizes rewards within sampled groups of trajectories. This reduces variance in policy updates, stabilizes training, and improves convergence efficiency compared to standard Proximal Policy Optimization (PPO). Unlike Flow-Steer’s multi-turn interaction with execution feedback, GRPO generates the entire workflow in one shot, limiting its ability to adapt based on intermediate results.
@@ E.3 Search-Based Workflow Methods
$\bullet$ AFlow (Zhang et al., 2024b): A workflow optimization framework that uses Monte Carlo Tree Search (MCTS) to explore the workflow space. The method systematically searches over predefined operator combinations through tree expansion and backpropagation, evaluating candidate workflows by execution outcomes. While effective at finding good operator sequences, AFlow lacks end-to-end learning capability and cannot learn from accumulated experience across different problems. Its search process can be computationally expensive as the workflow complexity increases.
@@ E.4 Agent with Reinforcement Learning Methods
$\bullet$ Router-R1: A router-style architecture where a small policy model learns to route queries to different processing paths using GRPO for policy optimization. The router makes a single decision per query without iterative refinement, selecting from a predefined set of workflow templates. While this approach is computationally efficient, the single-shot routing mechanism cannot adapt to intermediate execution results or refine workflows based on partial feedback, limiting its performance on complex multi-step reasoning tasks.

### v2

@@ 1 Introduction
In recent years, a variety of powerful agentic systems have been applied to solve a wide range of human problems (Ruan et al., 2023; Shen et al., 2023; Hong et al., 2024), gradually moving beyond single-turn question answering (QA) toward executable end-to-end task completion. In this process, workflow orchestration has become a key bridge from task goals to reproducible execution: by organizing operators into an executable workflow graph, systems can complete complex tasks with improved controllability, debuggability, and reusability (Zeng et al., 2023; Qian et al., 2024), as shown in Figure 1. However, in practice, workflow construction still heavily relies on manual drag-and-drop and rule-based configuration (Hu et al., 2024), making it costly to transfer across new tasks, new operator libraries, new model backends, or different application domains.
@@
To address these issues, three main paradigms of workflow orchestration have emerged, as shown in Figure 2. First, static workflow selection retrieves pre-defined workflows from a library based on task similarity (Wang et al., 2024). Second, offline workflow generation trains policy models via supervised fine-tuning (SFT) or group relative policy optimization (GRPO) (Shao et al., 2024) to generate workflows. Third, automated workflow optimization—such as AFlow (Zhang et al., 2024), GPTSwarm (Zhuge et al., 2024), and LATS (Zhou et al., 2024)—combines search and execution feedback to iteratively improve workflow structures.
@@
However, these methods still face several challenges: (i) High manual/heuristic dependence—rules and templates require continual maintenance and are often tightly coupled to specific scenarios, limiting reuse and generalization (Schick et al., 2023; Wang et al., 2024); (ii) Operator/backend lock-in (path lock-in)—existing approaches tend to rely on fixed operators or a single strong large language model (LLM) backend, making it difficult for a policy model to compose operators in a plug-and-play manner, so performance and robustness drop sharply when the operator library or backend environment changes (Ong et al., 2024; Chen et al., 2023); (iii) Sparse and unstable learning signals—training with only terminal correctness rewards can lead to shortcut behaviors (e.g., premature termination and oversimplified graphs) and reward hacking, and may further suffer from intra-group advantage collapse, making long-horizon credit assignment unstable (Yu et al., 2024; DeepSeek-AI, 2025).
@@ 2 Related Work
Agent Workflows. Before the rise of LLM agents, workflow automation was largely rule- or template-driven, with operators and control flow specified by hand. In the era of LLM agents, agent workflows improve long-horizon reliability via a plan–act–feedback loop (Yao et al., 2023; Wang et al., 2023; Hong et al., 2024). Existing studies can be grouped into three lines: single-agent decision making, which models tool use as sequential decisions (Qin et al., 2023) or interleaved reasoning and acting (Yao et al., 2023); orchestration, which uses LLM controllers for tool/model routing (Ong et al., 2024) or constrained application programming interface (API) planning to ground intent (Wang et al., 2024); and multi-agent collaboration, which relies on standard operating procedures (SOPs)/roles (Hong et al., 2024) or cross-team orchestration (Wu et al., 2023). For sustained capability, agentic tool reasoning (Feng et al., 2025; Luo et al., 2025c) and reusable workflow memory further improve efficiency (Wang et al., 2024, 2023), supporting scalable automation in practice (Qian et al., 2024). In this paper, we propose FlowSteer, an end-to-end RL framework that learns workflow orchestration from executable canvas feedback.
Reinforcement Learning for Agents. With recent progress in RL for LLM-based agents, agent RL models interaction as a long-horizon Markov decision process (MDP) (Zhou et al., 2024). In hierarchical multi-turn RL, coupling value learning with token-level policy learning eases delayed credit assignment (Zhou et al., 2024; Shen et al., 2026). For tool-chain control, step-grained shaping (Yu et al., 2024) and outcome feedback improve tool selection and self-correction (Feng et al., 2025). When retrieval is treated as an action, search rollouts learn think-then-search behaviors (Jin et al., 2025; Luo et al., 2025a). Moreover, large-scale RL motivates GRPO-style objectives based on verifiable or group-relative signals (DeepSeek-AI, 2025; Shao et al., 2024), and recent work explores end-to-end RL for collaborative prompting (Liu et al., 2025) and knowledge graph reasoning (Shen et al., 2025a, b).
@@ 4.3 Outcome-Directed End-to-End Reinforcement Learning
which measures the ratio between current policy $\pi_{\theta}$ and behavior policy $\pi_{\theta_{\mathrm{old}}}$ used for trajectory sampling, and the normalized advantage $\hat{A}(\tau_{i})=(R(\tau_{i})-\mu_{\mathrm{src}})/(\sigma_{\mathrm{src}}+\epsilon)$ uses within-group statistics (Shao et al., 2024) $\mu_{\mathrm{src}}$, $\sigma_{\mathrm{src}}$ partitioned by data source. The mask $\mathrm{mask}_{t}^{(i)}\in\{0,1\}$ is set to 1 for policy-generated tokens and 0 for environment feedback, ensuring gradients only affect policy decisions; $|\tau_{i}|_{\mathrm{mask}}$ is the masked token count. The $\mathrm{clip}(\cdot)$ operator stabilizes updates (Schulman et al., 2017), and the KL term regularizes toward reference $\pi_{\mathrm{ref}}$ with strength $\beta$.
@@ 5.1 Experimental Setup
Datasets. We evaluate FlowSteer on six in-distribution (IID) benchmarks covering three task categories: (i) mathematical reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b)), (ii) question answering (HotPotQA (Yang et al., 2018), SQuAD v2 (Rajpurkar et al., 2018)), and (iii) code generation (MBPP (Austin et al., 2021), HumanEval (Chen et al., 2021)). To assess generalization, we further test on six out-of-distribution (OOD) benchmarks: TriviaQA (Joshi et al., 2017), NaturalQuestions (Kwiatkowski et al., 2019), MathQA (Amini et al., 2019), AIME 2025 (AMC, 2025), APPS (Hendrycks et al., 2021a), and DS-1000 (Lai et al., 2023). More details of the datasets are provided in Appendix D.
Baselines. We compare FlowSteer against the following baselines: (i) direct LLM baselines (Qwen3-8B (Qwen Team, 2025), GPT-4o-mini (OpenAI, 2024)), (ii) standard fine-tuning methods (SFT (Ouyang et al., 2022), GRPO (Shao et al., 2024)), (iii) workflow-based methods (AFlow (Zhang et al., 2024)), and (iv) agent with RL methods (AgentFlow, Router-R1 (Zhang et al., 2025), Orchestrator (Su et al., 2025)). More details are provided in Appendix E.
@@ 5.6 Comparison of RL Algorithms (RQ5)
As shown in Table 6, we compare three RL algorithms (DAPO (Yu et al., 2025), GRPO, and CWRPO) on six IID benchmarks under identical training settings. CWRPO achieves optimal results across all task types with more stable training dynamics. The pairwise comparison matrices in Figure 5(c–e) show consistent superiority of CWRPO across math, QA, and code tasks. Unlike DAPO and GRPO that primarily focus on answer correctness, CWRPO jointly optimizes structural diversity and task performance through diversity-constrained rewards with conditional release. This design enables the policy to learn not only how to solve problems but also how to orchestrate workflows effectively, achieving better trade-off between workflow quality and answer accuracy.
@@ Appendix D Dataset Details
$\bullet$ AIME 2025 (AMC, 2025): Problems from the 2025 American Invitational Mathematics Examination, representing challenging competition-level mathematics. AIME problems require creative problem-solving and deep mathematical insight, with answers being integers from 0 to 999.
@@ E.2 Fine-Tuning Baselines
$\bullet$ GRPO (Group Relative Policy Optimization) (Shao et al., 2024): Group Relative Policy Optimization is a reinforcement learning algorithm that normalizes rewards within sampled groups of trajectories. This reduces variance in policy updates, stabilizes training, and improves convergence efficiency compared to standard Proximal Policy Optimization (PPO). Unlike Flow-Steer’s multi-turn interaction with execution feedback, GRPO generates the entire workflow in one shot, limiting its ability to adapt based on intermediate results.
@@ E.3 Search-Based Workflow Methods
$\bullet$ AFlow (Zhang et al., 2024): A workflow optimization framework that uses Monte Carlo Tree Search (MCTS) to explore the workflow space. The method systematically searches over predefined operator combinations through tree expansion and backpropagation, evaluating candidate workflows by execution outcomes. While effective at finding good operator sequences, AFlow lacks end-to-end learning capability and cannot learn from accumulated experience across different problems. Its search process can be computationally expensive as the workflow complexity increases.
@@ E.4 Agent with Reinforcement Learning Methods
$\bullet$ Router-R1: A router-style architecture where a small policy model learns multi-round routing and aggregation of queries to different processing paths via reinforcement learning. The router iteratively selects from a predefined set of workflow templates with refinement across rounds. While this approach is computationally efficient, the routing mechanism has limited ability to adapt to intermediate execution results or refine workflows based on partial feedback, limiting its performance on complex multi-step reasoning tasks.